The relative fitness of three genotypes at a particular locus are W11=0.65, W12=0.75, W22=0.9. Four populations have the following set of allele frequencies. In which population will the change in allele frequencies (due to selection) be the most rapid? (Show work) CLEAR hand writing or typing

A.    p = 0.86 & q = 0.14

B.    p = 1.0 & q = 0.0

C.   p = 0.05 & q = 0.95

D. p = 0.49 & q = 0.51
Answer #1

Fourth population has the largest change in allelic frequency so in this population, the change is rapid.

0.86 P = 0.9 population WIL = 0.65 + w mean = 0.14 Wiz - 0.75 Wz2 fitness p² Will + 2pq Wiz t q²W 22 0.480747 0.1806 + 0.0176APCO Date: Page New frequency op- p²W, 1 + pq w,2 - 0.65 = 1 =q²W 22 + p q W12 w q 0.65 W Change 1-1=0 3rd 3 population p =